# In-Context Learning as Nonparametric Conditional Probability Estimation: Risk Bounds and Optimality

### v1

@@ 3.1 In-Context Learning
In-context learning (ICL) refers to the phenomenon where a model performs prediction by conditioning on a sequence of input-output examples, known as a prompt, without updating its parameters (Brown et al., 2020; Garg et al., 2022). A prompt consists of a sequence $(\mathbf{x}_{1},\mathbf{y}_{1},\ldots,\mathbf{x}_{N},\mathbf{y}_{N},\mathbf{x}_{N+1})$, where each $\mathbf{y}_{i}=f(\mathbf{x}_{i})$ for some unknown function $f$, and $\mathbf{x}_{N+1}$ is a query input. The goal is to output a prediction $\hat{\mathbf{y}}(\mathbf{x}_{N+1})$ such that $\hat{\mathbf{y}}(\mathbf{x}_{N+1})\approx f(\mathbf{x}_{N+1})$.
In ICL, each task is specified by a joint distribution $P$ over $(\mathbf{x},\mathbf{y})\in[0,1]^{p}\times[K]$, where $P$ is drawn from a family $\mathcal{P}$. For a given task, the model observes $N$ i.i.d. examples from $P$ as a prompt and is then required to predict the label $\mathbf{y}_{N+1}$ for a query input $\mathbf{x}_{N+1}$, where $(\mathbf{x}_{N+1},\mathbf{y}_{N+1})\sim P$ independently of the prompt.
For a given task, we draw $(\mathbf{x}_{1},\mathbf{y}_{1}),\ldots,(\mathbf{x}_{N+1},\mathbf{y}_{N+1})\overset{\text{i.i.d.}}{\sim}P$, where the first $N$ pairs form the prompt and $\mathbf{x}_{N+1}$ is the query. Following Bai et al. (2023), we encode the prompt and the query into a matrix $\mathbf{Z}\in\mathbb{R}^{(p+K)\times(N+1)}$ as
$\mathbf{Z}=\begin{bmatrix}\mathbf{X}&\mathbf{x}_{N+1}\\ \mathbf{Y}&\mathbf{o}_{K}\end{bmatrix}$,
where $\mathbf{X}=(\mathbf{x}_{1},\ldots,\mathbf{x}_{N})\in\mathbb{R}^{p\times N}$, $\mathbf{Y}=(\mathbf{y}_{1},\ldots,\mathbf{y}_{N})\in\mathbb{R}^{K\times N}$, and $\mathbf{o}_{K}\in\mathbb{R}^{K}$ is a zero vector placeholder for the label of the query input $\mathbf{x}_{N+1}$, which the model is expected to predict based on the prompt.
We consider a dataset of $T$ tasks, denoted by $\mathcal{D}_{T}=\{(\mathbf{Z}^{(t)},\mathbf{y}^{(t)}_{N+1})\}_{t=1}^{T}$, where each prompt matrix $\mathbf{Z}^{(t)}\in[0,1]^{(p+K)\times(N+1)}$ is flattened into a vector in $0,1]^{d}$ with $d:=(p+K)(N+1)$ for notational simplicity.
@@ 3.2 The Multiclass Classification Model
Consider task $t$ with prompt $\mathbf{Z}^{(t)}=(\mathbf{X}^{(t)},\mathbf{Y}^{(t)},\mathbf{x}_{N+1}^{(t)})$, where $(\mathbf{X}^{(t)},\mathbf{Y}^{(t)})$ denotes the set of $N$ in-context examples and $\mathbf{x}_{N+1}^{(t)}$ is the query input. Let $\mathbf{y}_{N+1}^{(t)}\in\{0,1\}^{K}$ be the one-hot encoded label of the query, with $y_{N+1,k}^{(t)}$ denoting its $k$-th component. The true conditional class probabilities are defined as
@@
$\mathbf{p}_{0}(\mathbf{Z}^{(t)}):=\left(p_{1}^{0}(\mathbf{Z}^{(t)}),\dots,p_{K}^{0}(\mathbf{Z}^{(t)})\right)\in\mathcal{S}^{K},$
where $\mathcal{S}^{K}$ denotes the $K$-dimensional probability simplex.
To estimate the conditional class probabilities, a standard approach in deep learning is to minimize the empirical negative log-likelihood via (stochastic) gradient descent. Let $\mathcal{D}_{T}=\{(\mathbf{Z}^{(t)},\mathbf{y}_{N+1}^{(t)})\}_{t=1}^{T}$ be a dataset of $T$ tasks, where each $\mathbf{Z}^{(t)}$ denotes the prompt input and $\mathbf{y}_{N+1}^{(t)}\in\{0,1\}^{K}$ is a one-hot encoded label vector. Given a conditional class probability vector $\mathbf{p}(\mathbf{Z})=(p_{1}(\mathbf{Z}),\dots,p_{K}(\mathbf{Z}))$, the likelihood over $\mathcal{D}_{T}$ is defined as $\mathcal{L}(\mathbf{p}\mid\mathcal{D}_{T})=\prod_{t=1}^{T}\prod_{k=1}^{K}\left(p_{k}(\mathbf{Z}^{(t)})\right)^{y_{N+1,k}^{(t)}}$.
Taking the negative logarithm yields the empirical negative log-likelihood (or cross-entropy loss):
@@
where the logarithm is applied element-wise and denotes the transpose. This loss measures the average discrepancy between the predicted and true class distributions across tasks. Minimizing it corresponds to empirical risk minimization under the cross-entropy loss. The resulting estimator $\hat{\mathbf{p}}\in\mathcal{F}$ is obtained by optimizing over a function class $\mathcal{F}$, typically parameterized by a deep neural network and constrained to produce valid probability vectors.
We aim to investigate a non-asymptotic bound of the expected excess risk of the estimator $\hat{\mathbf{p}}$, as measured by the expected cross-entropy loss:
$\mathbb{E}_{\mathcal{D}_{T},(\mathbf{Z},\mathbf{y}_{N+1})}\left[\mathbf{y}_{N+1}^{\top}\log\frac{\mathbf{p}_{0}(\mathbf{Z})}{\hat{\mathbf{p}}(\mathbf{Z})}\right]=\mathbb{E}_{\mathcal{D}_{T},\mathbf{Z}}\left[\mathbf{p}_{0}(\mathbf{Z})^{\top}\log\frac{\mathbf{p}_{0}(\mathbf{Z})}{\hat{\mathbf{p}}(\mathbf{Z})}\right]=\mathbb{E}_{\mathcal{D}_{T},\mathbf{Z}}\left[{\rm KL}(\mathbf{p}_{0}(\mathbf{Z})\|\hat{\mathbf{p}}(\mathbf{Z}))\right],$
where $(\mathbf{Z},\mathbf{y}_{N+1})$ is independent of the data set $\mathcal{D}_{T}$, $\mathbf{p}_{0}(\mathbf{Z})$ is the true conditional distribution, and $\hat{\mathbf{p}}(\mathbf{Z})$ is the predicted distribution, and ${\rm KL}(\mathbf{p}_{0}(\mathbf{Z})\|\hat{\mathbf{p}}(\mathbf{Z}))$ denotes the Kullback–Leibler (KL) divergence between the two distributions, conditional on $\mathbf{Z}$ and $(\mathcal{D}_{T},\mathbf{Z})$, respectively.
@@
$\mathrm{KL}_{B}(\mathbf{p}_{0}(\mathbf{Z})\,\|\,\hat{\mathbf{p}}(\mathbf{Z}))=\mathbf{p}_{0}(\mathbf{Z})^{\top}\left(B\wedge\log\frac{\mathbf{p}_{0}(\mathbf{Z})}{\hat{\mathbf{p}}(\mathbf{Z})}\right),$
where $a\wedge b=\min\{a,b\}$. Therefore, the expected excess risk of the estimator $\hat{\mathbf{p}}$, with respect to the expected cross-entropy loss, is defined as the expected truncated Kullback-Leibler divergence:
$R_{B}(\mathbf{p}_{0},\hat{\mathbf{p}}):=\mathbb{E}_{\mathcal{D}_{T},\mathbf{Z}}\left[\mathrm{KL}_{B}(\mathbf{p}_{0}(\mathbf{Z})\,\|\,\hat{\mathbf{p}}(\mathbf{Z}))\right]=\mathbb{E}_{\mathcal{D}_{T},\mathbf{Z}}\left[\mathbf{p}_{0}(\mathbf{Z})^{\top}\left(B\wedge\log\frac{\mathbf{p}_{0}(\mathbf{Z})}{\hat{\mathbf{p}}(\mathbf{Z})}\right)\right].$
@@
[…]
$R(\mathbf{p}_{0},\hat{\mathbf{p}})=\mathbb{E}_{\mathcal{D}_{T},\mathbf{Z}}\left[KL(\mathbf{p}_{0}(\mathbf{Z}),\hat{\mathbf{p}}(\mathbf{Z}))\right].$
Due to the non-convexity of the function class $\mathcal{F}$, we introduce a parameter to quantify the gap between the empirical risk achieved by an estimator $\hat{\mathbf{p}}$ and the global minimum within $\mathcal{F}$. Formally, this empirical risk gap is defined as:
$\Delta_{T}(\mathbf{p}_{0},\hat{\mathbf{p}}):=\mathbb{E}_{\mathcal{D}_{T}}\left[-\frac{1}{T}\sum_{t=1}^{T}\mathbf{y}_{N+1}^{(t)\,\top}\log(\hat{\mathbf{p}}(\mathbf{Z}^{(t)}))-\min_{\mathbf{p}\in\mathcal{F}}\left(-\frac{1}{T}\sum_{t=1}^{T}\mathbf{y}_{N+1}^{(t)\,\top}\log(\mathbf{p}(\mathbf{Z}^{(t)}))\right)\right].$ (1)
@@ 3.3 Model Classes: Transformers and MLPs
We begin by defining the model classes for our estimator, employing a transformer encoder for prompt representation and a subsequent MLP for processing the encoded output. This hybrid architecture—utilizing a transformer as a feature extractor with an MLP classifier atop—constitutes a prevalent design paradigm in practice. Transformers excel at capturing complex contextual dependencies through self-attention mechanisms, generating enriched latent representations of sequential data. The subsequent MLP leverages these high-level features to perform efficient classification or regression, benefiting from the transformer’s representational capacity while preserving computational tractability. Empirical studies across NLP benchmarks (e.g., BERT+MLP for GLUE tasks) demonstrate that such configurations offer an optimal balance: transformers extract transferable linguistic features, while lightweight MLP heads enable rapid task-specific adaptation without exhaustive retraining. The decoupled design further facilitates modular optimization, allowing domain-specific fine-tuning of the transformer alongside dedicated calibration of the classifier.
@@
Adapting the architecture proposed in previous theoretical work on transformer-based models Gurevych et al. (2022); Bai et al. (2023); Jiao et al. (2024), we adopt a transformer encoder to process the input prompt.
Let $\mathbf{Z}\in\mathbb{R}^{D\times(N+1)}$ denote the input prompt, where $D=p+K$ is the feature dimension and $N+1$ is the in-context length. A transformer network $\boldsymbol{\phi}$ maps $\mathbf{Z}$ to an output of the same dimension through a sequence of $M$ transformer blocks, each comprising a self-attention (SA) and a feed-forward (FF) sublayer. The architecture is defined as:
@@
$F^{(\mathrm{SA})}(\mathbf{Z})=\mathbf{Z}+\sum_{s=1}^{H}\mathbf{W}_{O,s}\left(\mathbf{W}_{V,s}\mathbf{Z}\right)\left[\left(\mathbf{W}_{K,s}\mathbf{Z}\right)^{\top}\left(\mathbf{W}_{Q,s}\mathbf{Z}\right)\right]\odot\sigma_{H}\left(\left(\mathbf{W}_{K,s}\mathbf{Z}\right)^{\top}\left(\mathbf{W}_{Q,s}\mathbf{Z}\right)\right),$
@@
where $d=(p+K)(N+1)$ and $\|\cdot\|_{0}$ counts the number of nonzero elements.
@@
where $\|\cdot\|_{\infty}$ is the element-wise maximum norm.
@@
Assume that the MLP class $\mathcal{F}_{\mathrm{id}}(L,W,S)$ is $L_{1}$-Lipschitz with respect to the $\ell_{\infty}$ norm, i.e.,
@@
This assumption simplifies our analysis. Although the Lipschitz constant $L_{1}$ of an MLP can be bounded in terms of its depth, width, and weight norms, computing its exact value is NP-hard (Virmaux and Scaman, 2018). Moreover, obtaining tight upper bounds remains challenging (Fazlyab et al., 2019; Latorre et al., 2020; Kim et al., 2021). Therefore, we do not attempt to estimate it explicitly and instead assume a uniform bound $L_{1}$.

### v2

@@ 3.1 In-Context Learning
In-context learning (ICL) refers to the phenomenon where a model performs prediction by conditioning on a sequence of input-output examples, known as a prompt, without updating its parameters (Brown et al., 2020; Garg et al., 2022). A prompt consists of $N$ labeled examples together with an unlabeled query input, i.e., $(\mathbf{x}_{1},\mathbf{y}_{1},\ldots,\mathbf{x}_{N},\mathbf{y}_{N},\mathbf{x}_{N+1})$, where each $\mathbf{y}_{i}=f(\mathbf{x}_{i})$ for some unknown function $f$. The goal is to output a prediction $\hat{\mathbf{y}}(\mathbf{x}_{N+1})$ for the query input $\mathbf{x}_{N+1}$ such that $\hat{\mathbf{y}}(\mathbf{x}_{N+1})\approx f(\mathbf{x}_{N+1})$.
In ICL for multiclass classification, each task is specified by a joint distribution $P$ over $(\mathbf{x},\mathbf{y})\in[0,1]^{p}\times\{0,1\}^{K}$, where $\mathbf{y}$ is a one-hot label vector. The distribution $P$ is drawn from a family of probability measures $\mathcal{P}$ and may vary across tasks. For a given task, the model observes $N$ i.i.d. examples $(\mathbf{x}_{1},\mathbf{y}_{1}),\ldots,(\mathbf{x}_{N},\mathbf{y}_{N})$ from $P$ and is then required to predict the label $\mathbf{y}_{N+1}$ for a query input $\mathbf{x}_{N+1}$, where $(\mathbf{x}_{N+1},\mathbf{y}_{N+1})\sim P$ independently of the first $N$ labeled examples. Following Bai et al. (2023), we encode the prompt $(\mathbf{x}_{1},\mathbf{y}_{1},\ldots,\mathbf{x}_{N},\mathbf{y}_{N},\mathbf{x}_{N+1})$ into a matrix $\mathbf{Z}\in\mathbb{R}^{(p+K)\times(N+1)}$:
$\mathbf{Z}=\begin{bmatrix}\mathbf{X}&\mathbf{x}_{N+1}\\ \mathbf{Y}&\mathbf{0}_{K}\end{bmatrix}$,
where $\mathbf{X}=(\mathbf{x}_{1},\ldots,\mathbf{x}_{N})\in\mathbb{R}^{p\times N}$, $\mathbf{Y}=(\mathbf{y}_{1},\ldots,\mathbf{y}_{N})\in\mathbb{R}^{K\times N}$, and $\mathbf{o}_{K}\in\mathbb{R}^{K}$ is a zero vector placeholder for the label of the query input $\mathbf{x}_{N+1}$, which the model is expected to predict based on the prompt. We consider a data set of $T$ tasks, denoted by $\mathcal{D}_{T}=\{(\mathbf{Z}^{(t)},\mathbf{y}^{(t)}_{N+1})\}_{t=1}^{T}$, where each prompt matrix $\mathbf{Z}^{(t)}\in[0,1]^{(p+K)\times(N+1)}$ is flattened into a vector in $0,1]^{d}$ with $d:=(p+K)(N+1)$ for notational simplicity.
@@ 3.2 The Multiclass Classification Model
Consider task $t$ with prompt $\mathbf{Z}^{(t)}=(\mathbf{X}^{(t)},\mathbf{Y}^{(t)},\mathbf{x}_{N+1}^{(t)})$, where $(\mathbf{X}^{(t)},\mathbf{Y}^{(t)})$ denotes the set of $N$ in-context examples and $\mathbf{x}_{N+1}^{(t)}$ is the query input. Let $\mathbf{y}_{N+1}^{(t)}\in\{0,1\}^{K}$ be the one-hot label vector of the query, with $y_{N+1,k}^{(t)}$ denoting its $k$-th component. The true conditional class probabilities are defined as
@@
$\mathbf{p}_{0}(\mathbf{Z}^{(t)}):=\left(p_{1}^{0}(\mathbf{Z}^{(t)}),\dots,p_{K}^{0}(\mathbf{Z}^{(t)})\right)^{\top}\in\mathcal{S}^{K},$
where $\mathcal{S}^{K}$ denotes the $K$-dimensional probability simplex and $\top$ denotes the transpose.
To estimate the conditional class probabilities, a standard approach in deep learning is to minimize the empirical negative log-likelihood via (stochastic) gradient descent. Let $\mathcal{D}_{T}=\{(\mathbf{Z}^{(t)},\mathbf{y}_{N+1}^{(t)})\}_{t=1}^{T}$ be a data set of $T$ tasks. Given a conditional class probability vector $\mathbf{p}(\mathbf{Z})=(p_{1}(\mathbf{Z}),\dots,p_{K}(\mathbf{Z}))^{\top}$, the likelihood over $\mathcal{D}_{T}$ is defined as $\mathcal{L}(\mathbf{p}\mid\mathcal{D}_{T})=\prod_{t=1}^{T}\prod_{k=1}^{K}\left(p_{k}(\mathbf{Z}^{(t)})\right)^{y_{N+1,k}^{(t)}}$. Taking the negative logarithm yields the empirical negative log-likelihood (or cross-entropy loss):
@@
where the logarithm is applied element-wise. This loss measures the average discrepancy between the predicted and true class distributions across tasks. Minimizing it corresponds to empirical risk minimization under the cross-entropy loss. The resulting estimator $\hat{\mathbf{p}}\in\mathcal{F}$ is obtained by optimizing over a function class $\mathcal{F}$, typically parameterized by a deep neural network and constrained to produce valid probability vectors.
We aim to investigate a non-asymptotic bound of the expected excess risk of the estimator $\widehat{\mathbf{p}}$, as measured by the expected cross-entropy loss:
[…]
$=$ $\mathbb{E}_{\mathcal{D}_{T},\mathbf{Z}}\left[\mathbf{p}_{0}(\mathbf{Z})^{\top}\log\frac{\mathbf{p}_{0}(\mathbf{Z})}{\widehat{\mathbf{p}}(\mathbf{Z})}\right]$
$=$ $\mathbb{E}_{\mathcal{D}_{T},\mathbf{Z}}\left[{\rm KL}(\mathbf{p}_{0}(\mathbf{Z})\|\widehat{\mathbf{p}}(\mathbf{Z}))\right],$
where $(\mathbf{Z},\mathbf{y}_{N+1})$ is independent of the data set $\mathcal{D}_{T}$, $\mathbf{p}_{0}(\mathbf{Z})$ is the true conditional distribution, and $\widehat{\mathbf{p}}(\mathbf{Z})$ is the predicted distribution, and ${\rm KL}(\mathbf{p}_{0}(\mathbf{Z})\|\widehat{\mathbf{p}}(\mathbf{Z}))$ denotes the Kullback-Leibler (KL) divergence between the two distributions $\mathbf{p}_{0}(\mathbf{Z})$ and $\widehat{\mathbf{p}}(\mathbf{Z})$, conditional on $\mathbf{Z}$ and $(\mathcal{D}_{T},\mathbf{Z})$, respectively.
@@
$\mathrm{KL}_{B}(\mathbf{p}_{0}(\mathbf{Z})\,\|\,\widehat{\mathbf{p}}(\mathbf{Z}))=\mathbf{p}_{0}(\mathbf{Z})^{\top}\left(B\wedge\log\frac{\mathbf{p}_{0}(\mathbf{Z})}{\widehat{\mathbf{p}}(\mathbf{Z})}\right),$
where $a\wedge b=\min\{a,b\}$. Therefore, the expected excess risk of the estimator $\hat{\mathbf{p}}$ is defined as the expected truncated Kullback-Leibler divergence:
$R_{B}(\mathbf{p}_{0},\widehat{\mathbf{p}}):=\mathbb{E}_{\mathcal{D}_{T},\mathbf{Z}}\left[\mathrm{KL}_{B}(\mathbf{p}_{0}(\mathbf{Z})\,\|\,\widehat{\mathbf{p}}(\mathbf{Z}))\right]=\mathbb{E}_{\mathcal{D}_{T},\mathbf{Z}}\left[\mathbf{p}_{0}(\mathbf{Z})^{\top}\left(B\wedge\log\frac{\mathbf{p}_{0}(\mathbf{Z})}{\widehat{\mathbf{p}}(\mathbf{Z})}\right)\right].$
@@
The practice of truncating the log-likelihood ratio to ensure boundedness has been well established, see, for example, Wong and Shen (1995) and Bos and Schmidt-Hieber (2022). Truncation at level $B>0$ regularizes the KL divergence by controlling the influence of near-zero predictions, thereby preventing instability in the risk.
Furthermore, due to the non-convexity of the function class $\mathcal{F}$, we introduce a parameter to quantify the gap between the expected empirical risk achieved by an estimator $\hat{\mathbf{p}}$ and the global minimum within $\mathcal{F}$. Formally, this expected empirical risk gap is defined as:
$\Delta_{T}(\mathbf{p}_{0},\widehat{\mathbf{p}}):=\mathbb{E}_{\mathcal{D}_{T}}\left[-\frac{1}{T}\sum_{t=1}^{T}\mathbf{y}_{N+1}^{(t)\,\top}\log(\widehat{\mathbf{p}}(\mathbf{Z}^{(t)}))-\min_{\mathbf{p}\in\mathcal{F}}\left(-\frac{1}{T}\sum_{t=1}^{T}\mathbf{y}_{N+1}^{(t)\,\top}\log(\mathbf{p}(\mathbf{Z}^{(t)}))\right)\right].$ (1)
@@ 3.3 Model Classes: Transformers and MLPs
We begin by defining the model classes for our estimator, employing a transformer encoder for prompt representation and a subsequent MLP for processing the encoded output. This hybrid architecture–utilizing a transformer as a feature extractor with an MLP classifier atop–constitutes a prevalent design paradigm in practice. Transformers excel at capturing complex contextual dependencies through self-attention mechanisms, generating enriched latent representations of sequential data. The subsequent MLP leverages these high-level features to perform efficient classification or regression, benefiting from the transformer’s representational capacity while preserving computational tractability. Empirical studies across NLP benchmarks (e.g., BERT+MLP for GLUE tasks) demonstrate that such configurations offer an optimal balance: transformers extract transferable linguistic features, while lightweight MLP heads enable rapid task-specific adaptation without exhaustive retraining. The decoupled design further facilitates modular optimization, allowing domain-specific fine-tuning of the transformer alongside dedicated calibration of the classifier.
@@
Adapting the architecture proposed in previous theoretical work on transformer-based models (Gurevych et al., 2022; Bai et al., 2023; Jiao et al., 2024), we adopt a transformer encoder to process the input prompt.
Let $\mathbf{Z}\in\mathbb{R}^{D\times(N+1)}$ denote the input prompt, where $D=p+K$ is the feature dimension and $N$ is the in-context length. A transformer network $\boldsymbol{\phi}$ maps $\mathbf{Z}$ to an output of the same dimension through a sequence of $M$ transformer blocks, each comprising a self-attention (SA) and a feed-forward (FF) sublayer. The architecture is defined as:
@@
$F^{(\mathrm{SA})}(\mathbf{Z})=\mathbf{Z}+\sum_{s=1}^{H}\mathbf{W}_{O,s}\left(\mathbf{W}_{V,s}\mathbf{Z}\right)\left(\left(\mathbf{W}_{K,s}\mathbf{Z}\right)^{\top}\left(\mathbf{W}_{Q,s}\mathbf{Z}\right)\right)\odot\sigma_{H}\left(\left(\mathbf{W}_{K,s}\mathbf{Z}\right)^{\top}\left(\mathbf{W}_{Q,s}\mathbf{Z}\right)\right),$
@@
where $d=(p+K)(N+1)$, $D=p+K$, and $\|\cdot\|_{0}$ counts the number of nonzero elements.
@@
where $\|\cdot\|_{\infty}$ is the entrywise max-norm.
@@
The MLP class $\mathcal{F}_{\mathrm{id}}(L,W,S)$ is assumed to be $L_{1}$-Lipschitz with respect to the entrywise max-norm $\|\cdot\|_{\infty}$, that is,
@@
This assumption simplifies our theoretical analysis. Although the Lipschitz constant $L_{1}$ of an MLP can be bounded in terms of its depth, width, and weight norms, computing its exact value is NP-hard (Virmaux and Scaman, 2018). Moreover, obtaining tight upper bounds remains challenging (Fazlyab et al., 2019; Latorre et al., 2020; Kim et al., 2021). Therefore, we do not attempt to estimate it explicitly and instead assume a uniform bound $L_{1}$.